Solve the following system by any method:
{2(x +y) +8 =0
6x + 4y =-28
If the equations of the system are dependent, or if a system is inconsistent, so indicate.
In those cases enter dependent or inconsistent. Otherwise, enter your answer as an ordered pair.

Answer by Nate(3500)

2(x +y) +8 =0
2x + 2y + 8 = 0
2x = -2y - 8
x = -y - 4
-------------------------------------------------------------
6x + 4y =-28
6(-y - 4) +4y= -28
-6y - 24 + 4y = -28
-2y = -4
y = 2
-------------------------------------------
2x + 2y + 8 = 0
2x + 4 + 8 = 0
2x = -12
x = -6
*perpendicular due to different slopes*
(-6,2)
(independent,consistent)